# Extended mean field control problems with constraints: The generalized Fritz-John conditions and Lagrangian method

### v1

@@ 3.3 FJ optimality condition for abstract problem (26)
Let $(\widehat{X},\widehat{\alpha})\in C=\mathbb{M}^{2}\times\mathscr{U}[0,T]$ be an optimal pair to the C-(E)MFC problem (22) (equivalent to the abstract formulation (26)). Then, there exist (nontrivial) multipliers $r_{0}\geq 0$, $\mu^{i}\in\mathcal{M}^{+}[0,T]$ for $i\in I_{1}$, $\eta^{j}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$ for $j\in I_{2}$ and $\lambda\in\mathbb{M}^{2}$ which are not all zeros such that
$\begin{cases}\displaystyle\int_{[0,T]}\Phi_{i}(\widehat{X},\widehat{\alpha})(t)\mu^{i}(\mathrm{d}t)=0,~{}~{}\forall i\in I_{1},\\[13.99995pt] \displaystyle\mathbb{E}\left[\int_{0}^{T}\Psi_{j}(\widehat{X},\widehat{\alpha})(t)\eta^{j}_{t}\mathrm{d}t\right]=0,~{}~{}\forall j\in I_{2},\\[13.99995pt] \displaystyle r_{0}+\sum_{i\in I_{1}}\|\mu^{i}\|_{\rm TV}+\sum_{j\in I_{2}}\|\eta^{j}\|_{\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}}+\|\lambda\|_{\mathbb{M}^{2}}=1,\\[13.99995pt] \displaystyle-\xi_{1}=r_{0}D_{X}J(\widehat{X},\widehat{\alpha})(\cdot)+\sum_{i\in I_{1}}\int_{[0,T]}D_{X}\Phi_{i}(\widehat{X},\widehat{\alpha})(\cdot)(t)\mu^{i}(\mathrm{d}t)\\[8.00003pt] \displaystyle\quad\qquad+\sum_{j\in I_{2}}\mathbb{E}\left[\int_{0}^{T}D_{X}\Psi_{j}(\widehat{X},\widehat{\alpha})(\cdot)(t)\eta^{j}_{t}\mathrm{d}t\right]-\left\langle\lambda,D_{X}F(\widehat{X},\widehat{\alpha})(\cdot)\right\rangle_{\mathbb{M}^{2}}~{}~{}\text{\rm in}~{}\mathbb{M}^{2},\\[13.99995pt] \displaystyle-\xi_{2}=r_{0}D_{\alpha}J(\widehat{X},\widehat{\alpha})(\cdot)+\sum_{i\in I_{1}}\int_{[0,T]}D_{\alpha}\Phi_{i}(\widehat{X},\widehat{\alpha})(\cdot)(t)\mu^{i}(\mathrm{d}t)\\[8.00003pt] \displaystyle\quad\qquad+\sum_{j\in I_{2}}\mathbb{E}\left[\int_{0}^{T}D_{\alpha}\Psi_{j}(\widehat{X},\widehat{\alpha})(\cdot)(t)\eta^{j}_{t}\mathrm{d}t\right]-\left\langle\lambda,D_{\alpha}F(\widehat{X},\widehat{\alpha})(\cdot)\right\rangle_{\mathbb{M}^{2}}~{}~{}\text{\rm in}~{}\mathbb{L}_{\mathbb{F}}^{2},\end{cases}$ (33)
@@
Building upon the FJ optimality condition for the abstract problem (26), we obtain the stochastic maximum principle (SMP) for our C-(E)MFC problems, which is another main result of this paper.
Theorem 3.11 (SMP for extended MFC problem with constraints).
Let $(\widehat{X},\widehat{\alpha})\in C:=\mathbb{M}^{2}\times\mathscr{U}[0,T]$ be an optimal solution to our C-(E)MFC problem (22). There exist some $r_{0}\geq 0$, $\mu^{i}\in\mathcal{M}^{+}[0,T]$ for $i\in I_{1}$ and $\eta^{j}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$ for $j\in I_{2}$ such that
@@
    for $j\in I_{2}$, since $\eta_{t}^{j}\geq 0$, $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-$\mathrm{a.s.}$ and $\eta^{j}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$, we can define the (continuous) increasing process $A^{j}=(A_{t}^{j})_{t\in[0,T]}$ by $A_{t}^{j}=\int_{0}^{t}\eta_{s}^{j}\mathrm{d}s$. Then, for $\mathbb{P}$-$\mathrm{a.s.}$ $\omega$, the process $A^{j}(\omega)$ only increases on the set $\{t\in[0,T];~{}\psi^{j}(t,\widehat{X}_{t}(\omega),\widehat{\alpha}_{t}(\omega))=0\}$, $\mathrm{d}t$-$\mathrm{a.s.}$,
@@
$\qquad\left.-\sum_{i\in I_{1}}\int_{t}^{T}D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})_{t}\mu^{i}(\mathrm{d}s)-\sum_{j\in I_{2}}\nabla_{u}\psi^{j}(t,\widehat{X}_{t},\widehat{\alpha}_{t})\eta_{t}^{j},u-\widehat{\alpha}_{t}\right\rangle\geq 0.$ (34)
Here, $D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})_{t}$ is the evaluation of the continuous function $D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})$ at time $t$ and $\widehat{\rho}_{t}:={\cal L}(\widehat{X}_{t},\widehat{\alpha}_{t})$ for $t\in[0,T]$. The Hamiltonian $H^{r_{0}}:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})\times\mathbb{R}^{l}\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}\mapsto\mathbb{R}$ is defined by
@@
$\quad+\int_{t}^{T}\mathbb{E}^{\prime}\left[\partial_{\mu}H^{r_{0}}(s,\widehat{X}_{s}^{\prime},\widehat{\rho}_{s},\widehat{\alpha}_{s}^{\prime},Y_{s},Z_{s})(\widehat{X}_{s},\widehat{\alpha}_{s})\right]\mathrm{d}s-\sum_{i\in I_{1}}\int_{[t,T]}\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})\mu^{i}(\mathrm{d}s)$
$\quad-\sum_{j\in I_{2}}\int_{t}^{T}\nabla_{x}\psi^{j}(s,\widehat{X}_{s},\widehat{\alpha}_{s})\mathrm{d}A_{s}^{j}-\int_{t}^{T}Z_{s}\mathrm{d}W_{s}$ (36)
with $\widehat{\mu}_{T}=\mathcal{L}(\widehat{X}_{T})$. Here, $(\widehat{X}^{\prime},\widehat{\alpha}^{\prime})$ is an independent copy of $(\widehat{X},\widehat{\alpha})$ defined on the probability space $(\Omega^{\prime},\mathcal{F}^{\prime},P^{\prime})$ constructed following Lemma 3.2.
@@
$\quad-\sum_{i\in I_{1}}\int_{[0,T]}\mathbb{E}[\langle\nabla_{x}\phi^{i}(t,\widehat{X}_{t},\widetilde{\widehat{\alpha}}^{t}),L_{t}\rangle]\mu^{i}(\mathrm{d}t)-\sum_{j\in I_{2}}\mathbb{E}\left[\int_{0}^{T}\langle\nabla_{x}\psi^{j}(t,\widehat{X}_{t},\widehat{\alpha}_{t}),L_{t}\rangle\eta^{j}_{t}\mathrm{d}t\right]$ (39)
@@
with $\widehat{\rho}_{t}=\mathcal{L}(\widehat{X}_{t},\widehat{\alpha}_{t})$ for $t\in[0,T]$ and $\widehat{\mu}_{T}=\mathcal{L}(\widehat{X}_{T})$. We then define an element $p=(p_{t})_{t\in[0,T]}$ in $\mathbb{M}^{2}$ as follows, for $t\in[0,T]$,
@@
$\quad-\sum_{i\in I_{1}}\int_{[t,T]}\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})\mu^{i}(\mathrm{d}s)-\sum_{j\in I_{2}}\int_{t}^{T}\nabla_{x}\psi^{j}(s,\widehat{X}_{s},\widehat{\alpha}_{s})\eta^{j}_{s}\mathrm{d}s$
@@
where $MN:=\sum_{i=1}^{r}M_{i}L_{i}\in\mathbb{R}^{n}$ for any $M=(M_{1},\ldots,M_{r})\in\mathbb{R}^{n\times n\times r}$ and $L=(L_{1},\ldots,L_{r})\in\mathbb{R}^{n\times r}$. Here, the integral $\int_{[t,T]}\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})\mu^{i}(\mathrm{d}s)$ is well-defined because $\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})$ is Borel measurable w.r.t. time $s$ and $\int_{[t,T]}\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})\mu^{i}(\mathrm{d}s)$ is finite $\mathbb{P}$-$\mathrm{a.s.}$, which results from $\mathbb{E}[\int_{[t,T]}|\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})|\mu^{i}(\mathrm{d}s)]<\infty$ for $i\in I_{1}$. In addition, $\mathbb{E}[\int_{0}^{T}\nabla_{x}\psi^{j}(t,\widehat{X}_{t},\widehat{\alpha}_{t})\eta_{t}^{j}\mathrm{d}t]<\infty$, and hence $\int_{t}^{T}\nabla_{x}\psi^{j}(s,\widehat{X}_{s},\widehat{\alpha}_{s})\eta_{s}^{j}\mathrm{d}s$ is $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-$\mathrm{a.s.}$ well-defined. Applying Itô’s formula, we can derive that
@@
$\quad=\mathbb{E}\left[\int_{0}^{T}\langle\lambda_{t}^{1}-p_{t},\mathrm{d}L_{t}\rangle+\langle p_{T},L_{T}\rangle-\langle p_{0},L_{0}\rangle-\int_{0}^{T}\langle\mathrm{d}p_{t},L_{t}\rangle-\langle p,L\rangle_{T}\right]$
@@
$\quad-\sum_{i\in I_{1}}\int_{[0,T]}\mathbb{E}[\langle\nabla_{x}\phi^{i}(t,\widehat{X}_{t},\widetilde{\widehat{\alpha}}^{t}),L_{t}\rangle]\mu^{i}(\mathrm{d}t)-\sum_{j\in I_{2}}\int_{0}^{T}\mathbb{E}\left[\langle\nabla_{x}\psi^{j}(t,\widehat{X}_{t},\widehat{\alpha}_{t})\eta_{t}^{j},L_{t}\rangle\right]\mathrm{d}t$
@@
$\quad-\sum_{i\in I_{1}}\int_{[0,T]}\mathbb{E}[\langle\nabla_{x}\phi^{i}(t,\widehat{X}_{t},\widetilde{\widehat{\alpha}}^{t}),L_{t}\rangle]\mu^{i}(\mathrm{d}t)-\sum_{j\in I_{2}}\int_{0}^{T}\mathbb{E}\left[\langle\nabla_{x}\psi^{j}(t,\widehat{X}_{t},\widehat{\alpha}_{t})\eta_{t}^{j},L_{t}\rangle\right]\mathrm{d}t,$
@@
$\quad-\sum_{i\in I_{1}}\int_{[t,T]}\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})\mu^{i}(\mathrm{d}s)-\sum_{j\in I_{2}}\int_{t}^{T}\nabla_{x}\psi^{j}(s,\widehat{X}_{s},\widehat{\alpha}_{s})\eta^{j}_{s}\mathrm{d}s$
@@
$\quad+\int_{0}^{T}\left\langle\mathbb{E}^{\prime}\left[\partial_{\nu}b(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t},\widehat{\alpha}_{t})^{\top}{\lambda_{t}^{1}}^{\prime}+\partial_{\nu}\sigma(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t},\widehat{\alpha}_{t})^{\top}{\lambda_{t}^{2}}^{\prime}\right],\alpha_{t}-\widehat{\alpha}_{t}\right\rangle\mathrm{d}t$
$\quad-\left.\sum_{i\in I_{1}}\int_{[0,T]}\int_{0}^{t}\left\langle D_{\gamma}\phi^{i}(t,\widehat{X}_{t},\widetilde{\widehat{\alpha}}^{t})_{s},\alpha_{s}-\widehat{\alpha}_{s}\right\rangle\mathrm{d}s\mu^{i}(\mathrm{d}t)-\sum_{j\in I_{2}}\int_{0}^{T}\left\langle\nabla_{u}\psi^{j}(t,\widehat{X}_{t},\widehat{\alpha}_{t})\eta^{j}_{t},\alpha_{t}-\widehat{\alpha}_{t}\right\rangle\mathrm{d}t\right]$

### v2

@@ 3.3 FJ optimality condition for abstract problem (19)
Let $(\widehat{X},\widehat{\alpha})\in C=\mathbb{M}^{2}\times\mathscr{U}[0,T]$ be an optimal pair to the C-MFC problem (16) (equivalent to the abstract formulation (19)). Then, there exist (nontrivial) multipliers $r_{0}\geq 0$, $\mu^{i}\in\mathcal{M}^{+}[0,T]$ for $i\in I_{1}$, $\eta^{j}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$ for $j\in I_{2}$ and $\lambda\in\mathbb{M}^{2}$ which are not all zeros such that
$\begin{cases}\displaystyle\int_{[0,T]}\Phi_{i}(\widehat{X},\widehat{\alpha})(t)\mu^{i}(\mathrm{d}t)=0,~{}\forall i\in I_{1};\quad\mathbb{E}\left[\int_{0}^{T}\Psi_{j}(\widehat{X},\widehat{\alpha})(t)\eta^{j}_{t}\mathrm{d}t\right]=0,~{}\forall j\in I_{2},\\[13.99995pt] \displaystyle r_{0}+\sum_{i\in I_{1}}\|\mu^{i}\|_{\rm TV}+\sum_{j\in I_{2}}\|\eta^{j}\|_{\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}}+\|\lambda\|_{\mathbb{M}^{2}}=1,\\[13.99995pt] \displaystyle-\xi_{1}=r_{0}D_{X}J(\widehat{X},\widehat{\alpha})(\cdot)+\sum_{i\in I_{1}}\int_{[0,T]}D_{X}\Phi_{i}(\widehat{X},\widehat{\alpha})(\cdot)(t)\mu^{i}(\mathrm{d}t)\\[8.00003pt] \displaystyle\quad\qquad+\sum_{j\in I_{2}}\mathbb{E}\left[\int_{0}^{T}D_{X}\Psi_{j}(\widehat{X},\widehat{\alpha})(\cdot)(t)\eta^{j}_{t}\mathrm{d}t\right]-\left\langle\lambda,D_{X}F(\widehat{X},\widehat{\alpha})(\cdot)\right\rangle_{\mathbb{M}^{2}}~{}~{}\text{\rm in}~{}\mathbb{M}^{2},\\[13.99995pt] \displaystyle-\xi_{2}=r_{0}D_{\alpha}J(\widehat{X},\widehat{\alpha})(\cdot)+\sum_{i\in I_{1}}\int_{[0,T]}D_{\alpha}\Phi_{i}(\widehat{X},\widehat{\alpha})(\cdot)(t)\mu^{i}(\mathrm{d}t)\\[8.00003pt] \displaystyle\quad\qquad+\sum_{j\in I_{2}}\mathbb{E}\left[\int_{0}^{T}D_{\alpha}\Psi_{j}(\widehat{X},\widehat{\alpha})(\cdot)(t)\eta^{j}_{t}\mathrm{d}t\right]-\left\langle\lambda,D_{\alpha}F(\widehat{X},\widehat{\alpha})(\cdot)\right\rangle_{\mathbb{M}^{2}}~{}~{}\text{\rm in}~{}\mathbb{L}_{\mathbb{F}}^{2},\end{cases}$ (25)
@@
Building upon the FJ optimality condition for the abstract problem (19), we can obtain the stochastic maximum principle (SMP) for our C-MFC problems, which is another main result of this paper.
Theorem 3.12 (SMP for C-MFC problems).
Let $(\widehat{X},\widehat{\alpha})\in C:=\mathbb{M}^{2}\times\mathscr{U}[0,T]$ be an optimal solution to our C-MFC problem (16). There exist some $r_{0}\geq 0$, $\mu^{i}\in\mathcal{M}^{+}[0,T]$ for $i\in I_{1}$ and $\eta^{j}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$ for $j\in I_{2}$ such that
@@
    the Radon measure $\mu^{i}$ is supported on $\{t\in[0,T];~{}\mathbb{E}[\phi^{i}(t,\widehat{X}_{t},\widehat{\mu}_{t},\widetilde{\widehat{\alpha}}^{t})]=0\}$ for all $i\in I_{1}$,
@@
    for $j\in I_{2}$, since $\eta_{t}^{j}\geq 0$, $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-$\mathrm{a.s.}$ and $\eta^{j}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$, we can define the (continuous) increasing process $A^{j}=(A_{t}^{j})_{t\in[0,T]}$ by $A_{t}^{j}=\int_{0}^{t}\eta_{s}^{j}\mathrm{d}s$. Then, for $\mathbb{P}$-$\mathrm{a.s.}$ $\omega$, the process $A^{j}(\omega)$ only increases on the set $\{t\in[0,T];~{}\psi^{j}(t,\widehat{X}_{t}(\omega),\widehat{\rho}_{t},\widehat{\alpha}_{t}(\omega))=0\}$, $\mathrm{d}t$-$\mathrm{a.s.}$,
@@
$\qquad-\left.\sum_{i\in I_{1}}\int_{[t,T]}D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widetilde{\widehat{\alpha}}^{s})_{t}\mu^{i}(\mathrm{d}s)-\sum_{j\in I_{2}}\nabla_{u}\psi^{j}(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})\eta_{t}^{j}\right.$
$\qquad-\left.\sum_{j\in J_{2}}\mathbb{E}^{\prime}\left[\partial_{\nu}\psi^{j}(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t},\widehat{\alpha}_{t})\right]\eta_{t}^{j},u-\widehat{\alpha}_{t}\right\rangle\geq 0,~{}\forall u\in U.$ (26)
Here, $D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widetilde{\widehat{\alpha}}^{s})_{t}$ is the evaluation of the continuous function $D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widetilde{\widehat{\alpha}}^{s})$ at time $t$, $\widehat{\mu}_{t}:=\mathcal{L}(\widehat{X}_{t})$ and $\widehat{\rho}_{t}:={\cal L}(\widehat{X}_{t},\widehat{\alpha}_{t})$ for $t\in[0,T]$. The Hamiltonian $H^{r_{0}}:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})\times\mathbb{R}^{l}\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}\mapsto\mathbb{R}$ is defined by
@@
$\quad+\int_{t}^{T}\mathbb{E}^{\prime}\left[\partial_{\mu}H^{r_{0}}(s,\widehat{X}_{s}^{\prime},\widehat{\rho}_{s},\widehat{\alpha}_{s}^{\prime},Y_{s},Z_{s})(\widehat{X}_{s},\widehat{\alpha}_{s})\right]\mathrm{d}s$
$\quad-\sum_{i\in I_{1}}\int_{[t,T]}\left(\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widetilde{\widehat{\alpha}}^{s})+\mathbb{E}^{\prime}\left[\partial_{\mu}\phi^{i}(s,\widehat{X}_{s}^{\prime},\widehat{\mu}_{s},\widetilde{\widehat{\alpha}}^{s})(\widehat{X}_{s})\right]\right)\mu^{i}(\mathrm{d}s)$
$\quad-\sum_{j\in I_{2}}\int_{t}^{T}\left(\nabla_{x}\psi^{j}(s,\widehat{X}_{s},\widehat{\rho}_{s},\widehat{\alpha}_{s})+\partial_{\mu}\psi^{j}(s,\widehat{X}_{s}^{\prime},\widehat{\rho}_{s},\widehat{\alpha}_{s}^{\prime})(\widehat{X}_{s},\widehat{\alpha}_{s})\right)\mathrm{d}A_{s}^{j}-\int_{t}^{T}Z_{s}\mathrm{d}W_{s}$ (28)
Here, $(\widehat{X}^{\prime},\widehat{\alpha}^{\prime})$ is an independent copy of $(\widehat{X},\widehat{\alpha})$ defined on the probability space $(\Omega^{\prime},\mathcal{F}^{\prime},P^{\prime})$ constructed following Lemma 3.2.
@@
$\quad-\sum_{i\in I_{1}}\int_{[0,T]}\mathbb{E}\left[\left\langle\nabla_{x}\phi^{i}(t,\widehat{X}_{t},\widehat{\mu}_{t},\widetilde{\widehat{\alpha}}^{t})+\mathbb{E}^{\prime}\left[\partial_{\mu}\phi^{i}(t,\widehat{X}_{t}^{\prime},\widehat{\mu}_{t},\widetilde{\widehat{\alpha}}^{t})(\widehat{X}_{t})\right],L_{t}\right\rangle\right]\mu^{i}(\mathrm{d}t)$
$\quad-\sum_{j\in I_{2}}\mathbb{E}\left[\int_{0}^{T}\left\langle\nabla_{x}\psi^{j}(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})+\mathbb{E}^{\prime}\left[\partial_{\mu}\psi^{j}(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t},\widehat{\alpha}_{t})\right],L_{t}\right\rangle\eta^{j}_{t}\mathrm{d}t\right]$ (31)
@@
with $\widehat{\rho}_{t}=\mathcal{L}(\widehat{X}_{t},\widehat{\alpha}_{t})$ and $\widehat{\mu}_{t}=\mathcal{L}(\widehat{X}_{t})$ for $t\in[0,T]$ We then define an element $p=(p_{t})_{t\in[0,T]}$ in $\mathbb{M}^{2}$ as follows, for $t\in[0,T]$,
@@
$\quad-\sum_{i\in I_{1}}\int_{[t,T]}\left(\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widetilde{\widehat{\alpha}}^{s})+\mathbb{E}^{\prime}\left[\partial_{\mu}\phi^{i}(s,\widehat{X}_{s}^{\prime},\widehat{\mu}_{s},\widetilde{\widehat{\alpha}}^{s})(\widehat{X}_{s})\right]\right)\mu^{i}(\mathrm{d}s)$
$\quad-\sum_{j\in I_{2}}\int_{t}^{T}\left(\nabla_{x}\psi^{j}(s,\widehat{X}_{s},\widehat{\rho}_{s},\widehat{\alpha}_{s})+\mathbb{E}^{\prime}\left[\partial_{\mu}\psi^{j}(s,\widehat{X}_{s}^{\prime},\widehat{\rho}_{s},\widehat{\alpha}_{s}^{\prime})(\widehat{X}_{s},\widehat{\alpha}_{s})\right]\right)\eta^{j}_{s}\mathrm{d}s$
@@
where $MN:=\sum_{i=1}^{r}M_{i}L_{i}\in\mathbb{R}^{n}$ for any $M=(M_{1},\ldots,M_{r})\in\mathbb{R}^{n\times n\times r}$ and $L=(L_{1},\ldots,L_{r})\in\mathbb{R}^{n\times r}$. Here, the integral $\int_{[t,T]}\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widetilde{\widehat{\alpha}}^{s})\mu^{i}(\mathrm{d}s)$ is well-defined because $\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widetilde{\widehat{\alpha}}^{s})$ is Borel measurable w.r.t. time $s$ and $\int_{[t,T]}\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widetilde{\widehat{\alpha}}^{s})\mu^{i}(\mathrm{d}s)$ is finite $\mathbb{P}$-$\mathrm{a.s.}$, which results from $\mathbb{E}[\int_{[t,T]}|\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widetilde{\widehat{\alpha}}^{s})|\mu^{i}(\mathrm{d}s)]<\infty$ for $i\in I_{1}$. In addition, $\mathbb{E}[\int_{0}^{T}\nabla_{x}\psi^{j}(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})\eta_{t}^{j}\mathrm{d}t]<\infty$, and hence $\int_{t}^{T}\nabla_{x}\psi^{j}(s,\widehat{X}_{s},\widehat{\rho}_{s},\widehat{\alpha}_{s})\eta_{s}^{j}\mathrm{d}s$ is $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-$\mathrm{a.s.}$ well-defined. Applying Itô’s formula, we can derive that
@@
$=\mathbb{E}\left[\int_{0}^{T}\langle\lambda_{t}^{1}-p_{t},\mathrm{d}L_{t}\rangle+\langle p_{T},L_{T}\rangle-\langle p_{0},L_{0}\rangle-\int_{0}^{T}\langle\mathrm{d}p_{t},L_{t}\rangle-\langle p,L\rangle_{T}\right]$
@@
$\quad-\sum_{i\in I_{1}}\int_{[0,T]}\mathbb{E}\left[\left\langle\nabla_{x}\phi^{i}(t,\widehat{X}_{t},\widehat{\mu}_{t},\widetilde{\widehat{\alpha}}^{t})+\mathbb{E}^{\prime}\left[\partial_{\mu}\phi^{i}(t,\widehat{X}_{t}^{\prime},\widehat{\mu}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t})\right],L_{t}\right\rangle\right]\mu^{i}(\mathrm{d}t)$
$\quad-\sum_{j\in I_{2}}\int_{0}^{T}\mathbb{E}\left[\left\langle\nabla_{x}\psi^{j}(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})\eta_{t}^{j}+\mathbb{E}^{\prime}\left[\partial_{\mu}\psi^{j}(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t},\widehat{\alpha}_{t})\right],L_{t}\right\rangle\right]\mathrm{d}t$
@@
$\quad-\sum_{i\in I_{1}}\int_{[0,T]}\mathbb{E}\left[\left\langle\nabla_{x}\phi^{i}(t,\widehat{X}_{t},\widehat{\mu}_{t},\widetilde{\widehat{\alpha}}^{t}),L_{t}\right\rangle+\left\langle\mathbb{E}^{\prime}\left[\partial_{\mu}\phi^{i}(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t}^{\prime})\right],L_{t}^{\prime}\right\rangle\right]\mu^{i}(\mathrm{d}t)$
$\quad-\sum_{j\in I_{2}}\int_{0}^{T}\mathbb{E}\left[\left\langle\nabla_{x}\psi^{j}(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})\eta_{t}^{j},L_{t}\right\rangle+\left\langle\mathbb{E}^{\prime}\left[\partial_{\mu}\psi^{j}(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t}^{\prime},\widehat{\alpha}_{t}^{\prime})\right],L_{t}^{\prime}\right\rangle\right]\mathrm{d}t,$
@@
$\quad-\sum_{i\in I_{1}}\int_{[t,T]}\left(\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widetilde{\widehat{\alpha}}^{s})+\mathbb{E}^{\prime}\left[\partial_{\mu}\phi^{i}(s,\widehat{X}_{s}^{\prime},\widehat{\mu}_{s},\widetilde{\widehat{\alpha}}^{s})(\widehat{X}_{s})\right]\right)\mu^{i}(\mathrm{d}s)$
$\quad-\sum_{j\in I_{2}}\int_{t}^{T}\left(\nabla_{x}\psi^{j}(s,\widehat{X}_{s},\widehat{\rho}_{s},\widehat{\alpha}_{s})+\mathbb{E}^{\prime}\left[\partial_{\mu}\psi^{j}(s,\widehat{X}_{s}^{\prime},\widehat{\rho}_{s},\widehat{\alpha}_{s}^{\prime})(\widehat{X}_{s},\widehat{\alpha}_{s})\right]\right)\eta^{j}_{s}\mathrm{d}s$
@@
$\quad+\int_{0}^{T}\left\langle\mathbb{E}^{\prime}\left[\partial_{\nu}b(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t},\widehat{\alpha}_{t})^{\top}{\lambda_{t}^{1}}^{\prime}+\partial_{\nu}\sigma(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t},\widehat{\alpha}_{t})^{\top}{\lambda_{t}^{2}}^{\prime}\right],\alpha_{t}-\widehat{\alpha}_{t}\right\rangle\mathrm{d}t$
$\quad-\sum_{i\in I_{1}}\int_{[0,T]}\int_{0}^{t}\left\langle D_{\gamma}\phi^{i}(t,\widehat{X}_{t},\widehat{\mu}_{t},\widetilde{\widehat{\alpha}}^{t})_{s},\alpha_{s}-\widehat{\alpha}_{s}\right\rangle\mathrm{d}s\mu^{i}(\mathrm{d}t)$
$\quad-\left.\sum_{j\in I_{2}}\int_{0}^{T}\left\langle\left(\nabla_{u}\psi^{j}(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})+\mathbb{E}^{\prime}\left[\partial_{\nu}\psi^{j}(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t},\widehat{\alpha}_{t})\right]\right)\eta^{j}_{t},\alpha_{t}-\widehat{\alpha}_{t}\right\rangle\mathrm{d}t\right]$